Calculate the volume of acetylene at NTP produced by 100g of CaC2with water.

Grade:11

2 Answers

Vikas TU
7 years ago
CaC2(s) + 2H2O(l) --------------->      Ca(OH)2  +  C2H2
moles of CaC2 = 100/64  therfore moles of actelyne is also =  100/64.
Volume at NTP would be = moles of acetylene * 22.4  hence.
Jay
6 years ago
CaC2+2H2o ---Ca(OH)2+C2H2. Means 1mole of cac2 gives 1mole of acetylene or 22.4 lit of gasNow 64g of CaC2 gives 22.4 lit. So 100g of CaC2 gives 22.4* 100/64 = 35 lit of gas ( here * used for multiplication